a circle with centre(-3,-5)has one end of its diameter at (5,-3) find other end of the diameter.
Answer by mananth(16946)

centre(-3,-5)has one end of its diameter at (5,-3) find other end of the diameter
center of the circle is the mid point of ends of diameter
apply mid point formula
let The end point be (x,y)
(x+5)/2 = -3
x+5=-6
x=-11
(y-3)/2=-5
y-3=-10
y=-7
(-11,-7)